OpenStudy (anonymous):

Which statement about this data is true? 54 55 51 54 56 Only the mean and mode are the same. Only the mean and median are the same. Only the median and mode are the same. The mean, median, and mode are all the same.

OpenStudy (anonymous):

median: is 54 mode is 54 mean: is 54 so D is your answer

OpenStudy (anonymous):

heat dude is incorrect only the mode and the mean are the same so the first one is the correct answer

OpenStudy (anonymous):

i say YOU are wrong, mode is 54 clearly seen. mean is 54: 51+54+54+55+56=270/5=54 median is 54: 51, 54, 54, 55, 56 ^^ you have to order them from least to highest, and it doesn't matter if the numbers are repeated, the middle number will always me the median. here it is 54. if you don't know how to find median http://www.mathsisfun.com/median.html http://www.purplemath.com/modules/meanmode.htm so you're wrong:/

OpenStudy (anonymous):

edit: what was your median?